calculate the mass in golf a golf ball traveling 95 mph with a de Broglie wavelength of 4.7x10^-34m
1
Expert's answer
2020-06-05T05:15:30-0400

h = 6,626*10−34J*s

"\lambda=h/mv=>m=h/\lambda v= 6,626*10^−34J*s/(4.7*10^-34m*1.6**95000m/3600s)=0.0334kg=33.4g"